OpenStudy (anonymous):

If g(x) = x^2 + 3, find g(4).

jimthompson5910 (jim_thompson5910):

Replace every x with 4 and evaluate g(x) = x^2 + 3 g(4) = (4)^2 + 3 g(4) = ??

OpenStudy (anonymous):

ah now I get it , thank you :3

jimthompson5910 (jim_thompson5910):

yw